# Passive Beam Shaping via Binary-Coded Apertures

### v2

@@ V-A Common “Inkwell” Geometry and Relation to the Array Model
Fig. 4(a) shows the bare 3D-printed inkwell base. The aperture is discretized into a uniform candidate lattice with pitch $d_{0}=2.5~\mathrm{mm}$. Each candidate location is a square inkwell with a $2.2\times 2.2~\mathrm{mm}$ opening and depth $\approx 0.4~\mathrm{mm}$, printed into a dielectric base of total thickness $\approx 0.8~\mathrm{mm}$ (approximately $0.4$ mm bottom floor plus $0.4$ mm wall layer). The well walls (i) confine conductive ink during deposition and (ii) enforce a repeatable metallized footprint across prototypes.
@@
TABLE I: Common CAD/STL and fabrication dimensions for the 3D-printed “inkwell” reflector platform (all designs).
@@
Center-to-center pitch $d_{0}=2.5$ mm in $x$ and $y$ ($\approx\lambda/2$ at $60$ GHz)
@@
Ground plane Continuous copper tape/sheet on rear face (high reflectivity at 60 GHz)
@@ V-C Stencil Layer for Repeatable Binary Metallization
For spatial-masking prototypes, we fabricate a matching 3D-printed stencil that enforces the designed ON/OFF pattern during conductive-ink deposition. The stencil (thickness $\approx 0.8~\mathrm{mm}$) contains square openings of $2.1\times 2.1~\mathrm{mm}$ (slightly smaller than the $2.2\times 2.2~\mathrm{mm}$ well openings) to reduce edge bleeding and improve pattern fidelity. The stencil is closed everywhere except at intended ON locations, so that the designed binary sequence is mapped into a repeatable metallization pattern. Fig. 5 shows representative CAD/STL renderings of the inkwell base and a corresponding stencil for a $(\theta_{I},\theta_{T})=(30^{\circ},-60^{\circ})$ design case.
Figure 5: Inkwell base and matching stencil (CAD/STL). (a) Base with all wells present. (b) Stencil used to metallize selected wells (ON locations). (c) Optional “matched-base” variant in which OFF locations are removed in the print; in this work, we use the base matched to the stencil base (c)+ stencil(b), but either (a)+(b) or (c)+(b) can be used.
@@ V-D Prototype Classes and Fabrication Workflow
Using the common printing-and-deposition workflow, we fabricate three prototype classes which correspond to the designs analyzed in Sections III and IV. Each prototype is realized by (i) 3D-printing the inkwell base (and, when applicable, a matching stencil), (ii) applying a continuous copper ground plane to the rear face, and (iii) depositing conductive paint into the selected inkwells. For mask-coded designs, the stencil confines deposition to the intended ON locations to improve pattern fidelity and repeatability.
@@
Fig. 6 shows fabricated prototypes for two representative steering configurations, $(\theta_{I},\theta_{T})=(30^{\circ},-60^{\circ})$ and $(\theta_{I},\theta_{T})=(45^{\circ},-10^{\circ})$, comparing spatial-mask and diffraction-order realizations on the same inkwell platform. These prototypes are used in Section VII to validate the proposed beam-control methods through $60$ GHz azimuthal sweeps.
(a) Mask, $(\theta_{I},\theta_{T})=(30^{\circ},-60^{\circ})$
@@
(c) Mask, $(\theta_{I},\theta_{T})=(45^{\circ},-10^{\circ})$
@@ VI Performance Criteria and Evaluation Metrics
VI-A Angular Pattern Metrics
For a fixed incidence angle $\theta_{I}$, the theoretical angular power pattern is
@@ VI-A Angular Pattern Metrics
VI-B System-Level Mapping and Energy Efficiency
To connect target-direction performance to a system-level metric, we map the target-direction SNR to an achievable rate using the Shannon formula
@@ VII Results and Discussion
This section reports theoretical and experimental results for the proposed fully passive, fabrication-coded reflectors. Section VII-A characterizes 1-bit ON/OFF masking and diffraction-order steering, with bipolar ($\pm 1$) and ideal continuous-phase reflectors included as reference baselines. Section VII-B then presents 60 GHz OTA measurements and discusses agreement with theory and deployment implications.
VII-A Theoretical Results
Unless otherwise stated, theory is evaluated at 60 GHz ($\lambda=5$ mm) using a 1-D array-factor model with $M=35$ and $d_{0}/\lambda=0.5$. We report $\gamma(\theta;\theta_{I})=|p(\theta,\theta_{I})|^{2}/M^{2}$.
VII-A1 Single-target synthesis: $\theta_{I}=45^{\circ}$, $\theta_{T}=-10^{\circ}$
@@
Fig. 7 compares the normalized gain $\gamma(\theta;\theta_{I})$ for the all-ON aperture, the proposed ON/OFF mask, the bipolar ($\pm 1$) mask, and an ideal continuous-phase reflector. The all-ON aperture exhibits a dominant specular lobe at $\theta_{\mathrm{spec}}=-\theta_{I}=-45^{\circ}$ and weak non-specular response. In contrast, the ON/OFF mask forms a deterministic non-specular lobe toward $\theta_{T}=-10^{\circ}$ by selecting lattice locations via the cosine-threshold rule.
@@
The non-specular target response of ON/OFF masking exhibits a $9.94$ dB loss relative to the ideal continuous-phase upper bound, consistent with Lemma 2, which lower-bounds the maximum achievable normalized target-lobe gain by $1/\pi^{2}$ (i.e., $10\log_{10}(1/\pi^{2})=-9.94$ dB). This loss is fundamentally due to amplitude-only control where an ideal surface can phase-align all $M$ contributions at the target. ON/OFF masking can only retain or suppress elements and cannot re-phase unfavorable contributions. At the specular direction, the ON/OFF pattern incurs a smaller loss (approximately $6.9$ dB) because the all-ON aperture is already phase-aligned and applying a non-specular mask primarily reduces the effective aperture through thinning and redistributes energy into other angles. In particular, Lemma 1 predicts an asymptotic activation ratio near $50\%$, contributing an order-of-$3$ dB reduction, with the remaining loss attributable to the non-uniform spatial weighting introduced to form the non-specular lobe. The bipolar ($\pm 1$) mask allows sign reversals and therefore more closely approximates the ideal coherent sum at the target, yielding an $\approx 3.9$ dB loss, consistent with 1-bit phase quantization trends reported in [29].
@@
VII-A2 Diffraction-order steering (unconstrained aperture)
@@
VII-A3 Multi-target synthesis: $\theta_{I}=30^{\circ}$, $\theta_{T}\in\{-7.8^{\circ},-60^{\circ}\}$
@@
In the ideal continuous-phase case, the reflected energy is intentionally split across two directions, and therefore the peak gain at each target is reduced relative to the single-target setting. The ON/OFF and bipolar curves exhibit the same trend, with per-target peaks reduced by power sharing and by the 1-bit constraint. Notably, the characteristic loss behavior observed in the single-target case is preserved. When compared to the ideal continuous-phase reference, the ON/OFF masking exhibits $\approx 10~\mathrm{dB}$ non-specular penalty relative to ideal continuous-phase control that is consistent with the $1/\pi^{2}$ bound in Lemma 2, while the specular component experiences a lowerl reduction due to thinning and energy redistribution (as discussed for Fig. 7).
@@
VII-A4 Aperture-matched comparison
Diffraction-order designs can exhibit high target-lobe peaks in the unconstrained setting since selecting a larger period $\delta$ increases the physical aperture when $M$ is held fixed. To enable a fair “apples-to-apples” comparison, Fig. 11 evaluates Diffraction-order and dense ON/OFF masking under a fixed physical footprint. In this aperture-matched setting, increasing $\delta$ does not enlarge the reflector; instead, it reduces the number of active cells or elements that fit within the fixed length, thereby lowering the coherent gain available to the desired diffraction order.
As shown in Fig. 11(a), for the single-target case $(\theta_{I},\theta_{T})=(45^{\circ},-10^{\circ})$, the dense ON/OFF mask on the $\lambda/2$ maintains a larger effective number of contributing cells and therefore sustains stronger target response than the aperture-matched Diffraction-order design. Fig. 11(b) shows the same trend in the multi-target setting where periodic steering becomes less efficient when the period increase forces a sparser effective aperture. This reduces coherent accumulation at each intended non-specular direction. These results indicate that, under fixed-size deployment constraints, dense ON/OFF masking is typically preferable. In contrast, Diffraction-order is attractive when aperture expansion is permissible due to its higher peak gains.
@@ VII-A4 Aperture-matched comparison
Figure 12: Target-direction gain variation across frequency for $M=35$ elements, $\theta_{I}=45^{\circ}$, and target $\theta_{T}=-10^{\circ}$, normalized to the 60 GHz design point: $10\log_{10}\!\big(G_{T}(f)/G_{T}(f_{0})\big)$ with $f_{0}=60$ GHz. The reflector geometry (mask/period) is held fixed while sweeping 55–65 GHz. For the difraction order curve, the inter-element period is selected without constraining the physical aperture size.
VII-A5 Target-direction gain variation across frequency.
Fig. 12 plots the target-direction gain change relative to the 60 GHz design point, $10\log_{10}\!\big(G_{T}(f)/G_{T}(f_{0})\big)$, for three fixed-geometry reflectors (ON/OFF mask, diffraction-order, and ideal continuous-phase). For all schemes, the physical reflector (mask/period) is held constant while the carrier is swept from 55–65 GHz, so any degradation is due to wideband mismatch. The phase progression across the aperture scales with $k(f)=2\pi/\lambda(f)$, and the design that is phase-aligned at $f_{0}$ becomes progressively misaligned as $f$ moves away from $f_{0}$.
The ideal configurable baseline exhibits the smallest gain loss across the band, showing a smooth, approximately symmetric roll-off about 60 GHz. The proposed ON/OFF mask follows the ideal trend closely, with only a modest additional loss toward the band edges. This indicates that the binary amplitude gating is relatively robust to moderate fractional bandwidth. Although the mask is designed at 60 GHz, its effective aperture still retains significant coherent addition toward the target over the entire 55–65 GHz range.
In contrast, the diffraction-order (periodic) design shows severe frequency sensitivity, with deep notches reaching tens of decibels below the 60 GHz reference at certain frequencies. This behavior is inherent to periodic/grating-lobe steering where the target beam is produced by satisfying a wavelength-dependent diffraction condition, $\delta(\sin\theta_{I}+\sin\theta_{T})/\lambda(f)=n$. When $\lambda(f)$ changes while the physical period $\delta$ is fixed, the designed order shifts away from $\theta_{T}$ causing abrupt target-gain losses These results highlight that diffraction-order steering can provide strong steering at the design frequency with very simple fabrication, but it is substantially less robust in wideband operation than the dense ON/OFF mask and the ideal continuous-phase upper bound.

### v3

@@ V-A Common “Inkwell” Geometry and Relation to the Array Model
Fig. 5(a) shows the bare 3D-printed inkwell base used in this paper. The aperture is discretized into a uniform candidate lattice with pitch $d_{0}=2.5~\mathrm{mm}$. Each candidate location is a square inkwell with a $2.2\times 2.2~\mathrm{mm}$ opening and depth $\approx 0.4~\mathrm{mm}$, printed into a dielectric base of total thickness $\approx 0.8~\mathrm{mm}$ (approximately $0.4$ mm bottom floor plus $0.4$ mm wall layer). The well walls (i) confine conductive ink during deposition and (ii) enforce a repeatable metallized footprint across prototypes.
@@
TABLE I: Geometry, material, and fabrication parameters of the 3D-printed “inkwell” reflector platform used in this paper.
@@
Center-to-center pitch $d_{0}=2.5$ mm in $x$ and $y$ ($\approx\lambda/2$ at $60.48$ GHz)
@@
Ground plane Continuous copper tape/sheet on rear face (high reflectivity at 60.48 GHz)
@@ V-C Stencil Layer for Repeatable Binary Metallization
For spatial-masking prototypes, we fabricate a matching 3D-printed stencil that enforces the designed ON/OFF pattern during conductive-ink deposition. The stencil (thickness $\approx 0.8~\mathrm{mm}$) contains square openings of $2.1\times 2.1~\mathrm{mm}$ (slightly smaller than the $2.2\times 2.2~\mathrm{mm}$ well openings) to reduce edge bleeding and improve pattern fidelity. The stencil is closed everywhere except at intended ON locations so that the designed binary sequence is mapped into a repeatable metallization pattern. Fig. 6 shows representative computer-aided design (CAD) and stereolithography (STL) renderings of the inkwell base and a corresponding stencil for a $(\theta_{I},\theta_{T})=(30^{\circ},-60^{\circ})$ design case.
Figure 6: Inkwell base and matching stencil (CAD/STL views). (a) Base with all wells present. (b) Stencil used to metallize selected wells (ON locations). (c) Optional “matched-base” variant in which OFF locations are removed in the print; in this work, we use the base matched to the stencil (c)+ stencil(b), but either (a)+(b) or (c)+(b) can be used.
@@ V-D Prototype Classes and Fabrication Workflow
Using the common printing-and-deposition workflow, we fabricate three prototype classes which correspond to the designs analyzed in Sections III and IV. Each prototype is realized by (i) 3D-printing the inkwell base (and a matching stencil), (ii) applying a continuous copper ground plane to the rear face, and (iii) depositing conductive paint into the selected inkwells. For mask-coded designs, the stencil confines deposition to the intended ON locations to improve pattern fidelity and repeatability.
@@
Fig. 7 shows fabricated prototypes for two representative steering configurations, $(\theta_{I},\theta_{T})=(30^{\circ},-60^{\circ})$ and $(\theta_{I},\theta_{T})=(45^{\circ},-10^{\circ})$, comparing spatial-mask and diffraction-order realizations on the same inkwell platform. These prototypes are used in Section VIII to validate the proposed beam-control methods through $60.48$ GHz azimuthal sweeps.
(a) ON/OFF, $(\theta_{I},\theta_{T})=(30^{\circ},-60^{\circ})$
@@
(c) ON/OFF, $(\theta_{I},\theta_{T})=(45^{\circ},-10^{\circ})$
@@ VI Performance Criteria and Evaluation Metrics
VII Performance Criteria and Evaluation Metrics
This section defines the metrics used to evaluate the fully passive, fabrication-coded reflectors developed in this paper. The proposed apertures are designed to generate deterministic non-specular departure lobes toward intended user directions. Their performance is quantified in terms of (i) target-direction response and (ii) system-level impact under reflector-side operational-power constraints.
VII-A Angular Pattern Metrics
For a fixed incidence angle $\theta_{I}$, the theoretical angular power pattern is defined as
@@ VII-A Angular Pattern Metrics
VII-B System-Level Mapping and Energy Efficiency
To connect target-direction performance to a system-level metric, we map the target-direction signal-to-noise ratio (SNR) to an achievable rate using the Shannon formula
@@ VIII Results and Discussion
This section presents both theoretical and experimental results for the proposed fully passive coded reflectors. Section VIII-A analyzes the performance of 1-bit ON/OFF masking and diffraction-order steering using the proposed aperture-level models. Section VIII-B then reports over the air measurements and discusses their agreement with theory, along with implications for practical deployment.
The 1-D array-factor expression in (9) is used to generate the azimuth-plane plots, whereas the planar-aperture model in (4) is used to generate the corresponding 2-D angular maps and 3-D patterns. For the ON/OFF design, the ideal target-direction phase profile is first formed using (14), then converted to a binary mask using the cosine-threshold rule in (16), and finally evaluated in the array model. The resulting binary vector $\mathbf{b}$ directly determines the metallized cell locations in the fabricated reflector, i.e., entries with $b_{m}=1$ identify the wells that are filled with conductive ink, while entries with $b_{m}=0$ remain unmetallized. Thus, the same analytically derived mask is used both to generate the theoretical ON/OFF plots and to define the metallization pattern of the 3D-printed prototype.
For the diffraction-order design, the required period is first obtained from the closed-form design rule in (31), and the resulting angular response is then computed using the uniform-period model in (27). That analytically determined period is also used to set the spacing of the active metallized columns in the fabricated 3D-printed diffraction-order prototype. In the 2-D evaluations, the same 1-D azimuthal design is replicated across rows to form the corresponding 2-D coded aperture. Accordingly, all theoretical plots reported in this section are analytical evaluations of the proposed aperture-level models, and the same analytical design parameters are used to define the physical metallization pattern and spacing of the fabricated prototypes.
To position the proposed approach relative to closely related designs, several benchmark cases are used throughout the theoretical and experimental evaluation. In the theoretical results, the proposed ON/OFF reflector is compared against three reference designs: a 1-bit ($\pm 1$) benchmark motivated by discrete-state RIS literature [29], a conventional all-ON passive reflector baseline representing non-engineered specular reflection as in prior passive mmWave reflector studies [19], and an ideal continuous-phase reflector serving as an upper-bound reference for target-direction gain, consistent with common RIS/LIS benchmarking practice [6, 7]. In the experimental results, the fabricated designs are compared against both the passive all-ON reflector and a smooth copper plate of the same aperture size. The copper plate serves as a conductor reference to verify that the all-ON metallized inkwell reflector exhibits comparable specular behavior. These benchmarks help clarify that the contribution of this work is a fully passive, fabrication-friendly coded-aperture reflector framework that enables target-direction beam shaping without requiring active tuning circuitry or a new 3-D geometry redesign for each beam objective.
VIII-A Theoretical Results
Unless otherwise stated, all theoretical results and figures in this section are obtained by evaluating the proposed analytical aperture-level models in MATLAB. No full-wave unit-cell simulations are used in this work. The ON/OFF results are generated using the formulation in Section III, whereas the diffraction-order results are generated using the periodic steering formulation in Section IV. Unless otherwise stated, the analysis is carried out at 60 GHz using $M=35$ and $d_{0}/\lambda=0.5$.
VIII-A1 Single-target synthesis: $\theta_{I}=45^{\circ}$, $\theta_{T}=-10^{\circ}$
@@
Fig. 8 compares the normalized gain $\gamma(\theta;\theta_{I})$ for four reflector configurations: the conventional all-ON passive aperture, the proposed ON/OFF mask, a bipolar ($\pm 1$) 1-bit reflector included as a discrete-state benchmark [29], and an ideal continuous-phase reflector included as an upper-bound reference [6, 7]. The all-ON aperture exhibits the expected specular-dominated response, with its strongest lobe at $\theta_{\mathrm{spec}}=-\theta_{I}=-45^{\circ}$ and only weak non-specular reflection. By contrast, the proposed ON/OFF mask produces a deterministic non-specular lobe toward $\theta_{T}=-10^{\circ}$ by activating a subset of lattice locations according to the cosine-threshold rule.
@@
The non-specular target response of ON/OFF masking exhibits a $9.94$ dB loss relative to the ideal continuous-phase upper bound. This is consistent with Lemma 2, which lower-bounds the maximum achievable normalized target-lobe gain by $1/\pi^{2}$ (i.e., $10\log_{10}(1/\pi^{2})=-9.94$ dB). This loss is fundamentally due to amplitude-only control where an ideal surface can phase-align all $M$ contributions at the target. ON/OFF masking can only retain or suppress elements and cannot re-phase unfavorable contributions. At the specular direction, the ON/OFF pattern incurs a smaller loss (approximately $6.9$ dB) because the all-ON aperture is already phase-aligned and applying a non-specular mask primarily reduces the effective aperture through thinning and redistributes energy into other angles. In particular, Lemma 1 shows an asymptotic activation ratio near $50\%$, contributing an order-of-$3$ dB reduction, with the remaining loss attributable to the non-uniform spatial weighting introduced to form the non-specular lobe. The bipolar ($\pm 1$) mask allows sign reversals and therefore more closely approximates the ideal coherent sum at the target, yielding an $\approx 3.9$ dB loss, consistent with 1-bit phase quantization trends reported in [29].
@@
VIII-A2 Diffraction-order steering (unconstrained aperture)
@@
VIII-A3 Multi-target synthesis: $\theta_{I}=30^{\circ}$, $\theta_{T}\in\{-7.8^{\circ},-60^{\circ}\}$
@@
In the ideal continuous-phase case, the reflected energy is intentionally split across two directions, and therefore the peak gain at each target is reduced relative to the single-target setting. The ON/OFF and bipolar curves exhibit the same trend, with per-target peaks reduced by power sharing and by the 1-bit constraint. Notably, the characteristic loss behavior observed in the single-target case is preserved. When compared to the ideal continuous-phase reference, the ON/OFF masking exhibits $\approx 10~\mathrm{dB}$ non-specular penalty relative to ideal continuous-phase control, while the specular component experiences a lower reduction due to thinning and energy redistribution (as discussed for Fig. 8).
@@
VIII-A4 Aperture-matched comparison
Diffraction-order designs can exhibit high target-lobe peaks in the unconstrained setting since selecting a larger period $\delta$ increases the physical aperture when $M$ is held fixed. To enable a fair “apples-to-apples” comparison, Fig. 12 evaluates diffraction-order and dense ON/OFF masking under a fixed physical footprint. In this aperture-matched setting, increasing $\delta$ reduces the number of active cells or elements that fit within the fixed length and thereby lowers the coherent gain to the desired diffraction order.
As shown in Fig. 12(a), for the single-target case $(\theta_{I},\theta_{T})=(45^{\circ},-10^{\circ})$, the dense ON/OFF mask on the $\lambda/2$ maintains a larger effective number of contributing cells and therefore sustains stronger target response than the aperture-matched diffraction-order design. Fig. 12(b) shows the same trend in the multi-target setting where periodic steering becomes less efficient due to a sparser effective aperture. This reduces coherent accumulation at each intended non-specular direction.
The associated sidelobe behavior follows from the underlying aperture structures. In the ON/OFF case, binary amplitude quantization introduces additional spatial harmonics and does not fully suppress the residual specular component, leading to stronger secondary lobes than in the ideal continuous-phase case. In the diffraction-order case, additional lobes are an inherent consequence of the periodic aperture structure, since the desired beam is realized as a selected diffraction order of a passive periodic reflector. Overall, these results indicate that under fixed-size deployment constraints, dense ON/OFF masking is typically preferable because it preserves a larger effective aperture and stronger target-direction gain. By contrast, diffraction-order steering is attractive when aperture expansion is permitted since its synthesis is simple and analytically tractable.
@@ VIII-A4 Aperture-matched comparison
VIII-A5 Two-Dimensional Angular Response and Secondary-Lobe Behavior
To complement the azimuth-plane cuts, Figs. 1316 show representative 3-D angular patterns and 2-D angular response maps obtained from the planar-aperture model in (4) for both the single-beam and multi-beam cases using a $35\times 35$ lattice. These results are consistent with the 1-D azimuthal analysis, while also revealing the response variation with the second angular coordinate $\phi$. Since the synthesis in this paper is carried out for azimuth-plane steering, the strongest response remains centered near $\phi=0$. As $|\phi|$ increases, the response decreases and broadens because the design does not impose independent elevation steering.
For the single-beam case, Figs. 13 and 14 correspond to $\theta_{I}=45^{\circ}$ and $\theta_{T}=-10^{\circ}$. The passive all-ON reflector exhibits the expected specular response with dominant energy near $\theta_{\mathrm{spec}}=-45^{\circ}$. The ideal fixed-phase reflector concentrates energy near the desired target direction. The proposed ON/OFF design also forms a deterministic non-specular lobe toward $\theta_{T}=-10^{\circ}$, but with broader secondary structure due to binary quantization and residual specular leakage. The diffraction-order reflector likewise produces a lobe near the target direction, but its periodic structure leads to more pronounced secondary lobes.
For the multi-beam case, Figs. 15 and 16 correspond to $\theta_{I}=30^{\circ}$ with target directions $\theta_{T}\in\{-7.8^{\circ},-60^{\circ}\}$. The passive all-ON reflector again remains predominantly specular, whereas the ideal fixed-phase reflector produces two intended lobes near the desired target directions. The ON/OFF reflector also enhances both target directions, confirming that deterministic multi-beam shaping can be approximated with a fully passive binary aperture. The diffraction-order reflector shown in these figures is a single-target periodic reference and therefore does not realize both target beams simultaneously. Overall, the 2-D aperture results reinforce the same main conclusion as the 1-D analysis. Fully passive binary-coded apertures can realize deterministic non-specular beam shaping on a fixed dense lattice, while diffraction-order steering offers simpler periodic synthesis at the cost of stronger secondary lobes and reduced flexibility for multi-beam operation.
Figure 13: Three-dimensional angular response patterns at 60 GHz for the representative single-beam case $\theta_{I}=45^{\circ}$ and target $\theta_{T}=-10^{\circ}$. The responses are normalized to the peak of each design at 60 GHz.
Figure 14: Two-dimensional angular response maps at 60 GHz for the representative single-beam case $\theta_{I}=45^{\circ}$ and target $\theta_{T}=-10^{\circ}$. The plots show the normalized response of the proposed ON/OFF reflector, the diffraction-order reflector, the passive all-ON reflector, and the ideal fixed-phase reflector, evaluated using the 2-D aperture model.
Figure 15: Three-dimensional angular response patterns at 60 GHz for the representative multi-beam case $\theta_{I}=30^{\circ}$ with target directions $\theta_{T}\in\{-7.8^{\circ},-60^{\circ}\}$.
Figure 16: Two-dimensional angular response maps at 60 GHz for the representative multi-beam case $\theta_{I}=30^{\circ}$ with target directions $\theta_{T}\in\{-7.8^{\circ},-60^{\circ}\}$. The plots show the normalized response of the proposed ON/OFF reflector, a diffraction-order reflector, the passive all-ON reflector, and the ideal fixed-phase multi-beam reflector, evaluated using the 2-D aperture model.
VIII-A6 Wideband Sensitivity of Fixed Passive Designs
Figs. 17 and 18 quantify the wideband sensitivity of the passive designs for the single-beam and multi-beam cases. In both cases, the responses are normalized to the corresponding 60 GHz design-point value so that the off-design loss can be observed directly as the carrier frequency is varied while the physical reflector geometry is held fixed.
For the single-beam case in Fig. 17, all designs exhibit reduced target-direction response as the operating frequency moves away from the design point. The ON/OFF reflector shows a more gradual reduction in target-direction response than the diffraction-order design. This behavior is expected because diffraction-order steering depends explicitly on the ratio between the fixed spatial period and the wavelength, so the selected diffraction order shifts more rapidly as frequency changes. By contrast, the dense ON/OFF aperture is comparatively more tolerant to moderate off-design operation, although it also experiences loss away from the design frequency.
A similar trend is observed in the multi-beam case shown in Fig. 18. The results again show that fixed passive beam patterns are frequency-sensitive, with the diffraction-order design exhibiting stronger degradation away from the 60 GHz design point. Although the ON/OFF reflector is not frequency-invariant, it maintains more stable target-direction behavior over moderate frequency variation and therefore provides a more robust passive alternative under fixed-aperture operation.
Figure 17: Target-direction gain variation versus frequency for the single-beam case $\theta_{I}=45^{\circ}$ and $\theta_{T}=-10^{\circ}$, normalized to the corresponding 60 GHz response of each design.
Figure 18: Target-direction gain variation versus frequency for the multi-beam case $\theta_{I}=30^{\circ}$ with target directions $\theta_{T}=-7.8^{\circ}$ and $\theta_{T}=-60^{\circ}$, normalized to the corresponding 60 GHz response of each design.